SHIFT CHECK LIST (To print more copies go to the wiki, look under ”Run Plans”)

DATE (month-day-year) Time (24 hour clock) Shift Checklist Recorder

Beam Energy (overview) (MeV) Beam Current (overview)(nA) IPM3H00AX/AY (medm) IPM3H00BX/BY (medm) Slow Raster Circular? Fast Raster Rmax,Ix,Iy (medm)

Run Number (recent one!) HMS Trigger Rate (scalers) BETA1 Trigger Rate (scalers) BETA2 Trigger Rate (scalers) Computer Deadtime

Electronics Room: WC Thresholds HMS BCM UP Temp. ok?

GAS SHED KEY is #29 Argon Bottle High/Low (PSI) Ethane Bottle High/Low (PSI) Helium Bottle High (PSI) Helium Flow (CFM) Mixer Flow Ar/Eth Mixer outlet pressure Alcohol Temp (degC) Controller Temp (degC) Alcohol Bubbling? Recorded in Gas Logbook?

Column  1  2  3  4  5  6

PAGE 1 OF 2 DATE (month-day-year) Time (24 hour clock) Shift Checklist Recorder

Column  1  2  3  4  5  6

HMS Slits (LAT HCC18) Scattering Chamber Vacuum Target HMS Shutter Open? (TV) HMS angle (TV/GUI) Hall Survey (TV) HMS Cer. Temp/Press (TV)

HMS HV (OK?) BETA HV (OK?) Bigcal HV (OK?)

HMS Q1 Current polarity/setpt HMS Q1 Hall Field HMS Q2 Current polarity/setpt HMS Q2 Hall Field HMS Q3 Current polarity/setpt HMS Q3 Hall Field HMS D Current polar./readback HMS D NMR Field HMS Q1 N2/He Level HMS Q2 N2/He Level HMS Q3 N2/He Level HMS D N2/He Level

hclog the following screens: BPM medm screen

scan into hclog when page is full Check List Complete!

PAGE 2 OF 2

(To edit: /group/sane/www/general/SANE_check_list.tex
then dvips -tletter -Ppdf SANE_check_list; ps2pdf SANE_check_list.ps).